OpenStudy (anonymous):

Simplify:

OpenStudy (anonymous):

\[\frac{ 18y^3 }{ 36xy^2 }\]

OpenStudy (anonymous):

Choices: 1/2x y/2x y^2/ 2x 1/ 2xy

jimthompson5910 (jim_thompson5910):

Reduce 18/36 to get ______

OpenStudy (anonymous):

I tried to simplify by 18y^2 but it doesn't work.

OpenStudy (anonymous):

and by 18y.

jimthompson5910 (jim_thompson5910):

break it up like this \[\large \frac{ 18y^3 }{ 36xy^2 }\] \[\large \frac{ 18*1*y^3 }{ 36*x*y^2 }\] \[\large \frac{18}{36}*\frac{1}{x}*\frac{y^3}{y^2}\]

jimthompson5910 (jim_thompson5910):

then reduce/simplify each piece

OpenStudy (anonymous):

1 - x y 2

jimthompson5910 (jim_thompson5910):

1/x stays as 1/x

jimthompson5910 (jim_thompson5910):

y^3 divided by y^2 is y^(3-2) = y^1 = y

jimthompson5910 (jim_thompson5910):

\[\large \frac{ 18y^3 }{ 36xy^2 }\] \[\large \frac{ 18*1*y^3 }{ 36*x*y^2 }\] \[\large \frac{18}{36}*\frac{1}{x}*\frac{y^3}{y^2}\] \[\large \frac{1}{2}*\frac{1}{x}*\frac{y}{1}\] \[\large \frac{1*1*y}{2*x*1}\] \[\large \frac{y}{2x}\]

OpenStudy (anonymous):

Ohhh okay, thank you(:

jimthompson5910 (jim_thompson5910):

yw